OpenStudy (anonymous):

which number(s) cannot be used in the following expressions: a^2-36/a^2+9a+36 a.-3 b.3 c.-6 d.6 e.18 f.-18

OpenStudy (anonymous):

Numbers that "cannot" be used are ones where the denominator is zero. In other words, the values of \(a\) such that \(a^2+9a+36=0\).

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

Do you agree with me that fractions can't have zero as a denominator?

OpenStudy (anonymous):

yes i know that part , so how do i figure out what numbers i cant use?

OpenStudy (anonymous):

Do you agree that \(\dfrac{a^2-36}{a^2+9a+36}\) is a fraction with denominator \(a^2+9a+36\)?

OpenStudy (anonymous):

If so, the values of \(a\) that you cannot use are the ones that make the denominator \(a^2+9a+36\) zero, which are the solutions to \(a^2+9a+36=0\), which is what I mentioned above.

OpenStudy (anonymous):

whats the next step?

OpenStudy (anonymous):

Solve \(a^2+9a+36=0\).

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

Start by factoring the left-hand side.

OpenStudy (anonymous):

Or, use the quadratic formula if you are really stuck.

OpenStudy (anonymous):

i tried to use the quadratic formula i didnt work

OpenStudy (anonymous):

The quadratic formula always works, @tbrooks3. ;-) What did you get?

OpenStudy (anonymous):

Ah I see...you probably got a negative under the square root, right? That means there are no real solutions, which means that the given expression can take in ANY number without restriction.

OpenStudy (anonymous):

so a ,c,f ?

OpenStudy (anonymous):

What? No! None of the answers are correct. There are NO restrictions.

OpenStudy (anonymous):

what does that mean?

OpenStudy (anonymous):

None of the answers given are correct. There is no number \(a\) could be that is not allowed.

OpenStudy (anonymous):

but i have to pick one of the multiple choice answers

OpenStudy (anonymous):

If you don't believe me, look here to see that the domain is all real numbers. http://www.wolframalpha.com/input/?i=(x%5E2-36)%2F(x%5E2%2B9x%2B36) Either the problem is put in wrong, or you copied it wrong. The function you gave doesn't have an answer for the question given.

OpenStudy (anonymous):

ohhh your right sorry it should say +18 on the denominator not +36 @yakeyglee

OpenStudy (anonymous):

*sigh*...okay so what are the solutions to \(a^2+9a+18=0\)?

OpenStudy (anonymous):

i still have no clue. i have a quadratic calculator but its not working right now

OpenStudy (anonymous):

...okay, so do it by hand. That's an easy one to factor, anyway. Give it a bit of thought...what numbers sum to 9 and multiply to 18?

OpenStudy (anonymous):

6,3

OpenStudy (anonymous):

Good...so the factored form is \(a^2+9a+18=(a+3)(a+6)=0\).

OpenStudy (anonymous):

?

OpenStudy (anonymous):

b & c ?